OpenStudy (anonymous):

Gotta question for somone me and my family are diagreeing on.... ok so what is 3x-3=3/4x(2x+12)?

OpenStudy (anonymous):

(3x-3)(8x^2+48x)=3

OpenStudy (anonymous):

kay keep going?

OpenStudy (anonymous):

24x^3+144x^2-24x^2-144x=3

OpenStudy (anonymous):

24x^3+120x^2-144x=3 8x^3+40x^2-48x=1

OpenStudy (anonymous):

Kay whats me anwser?

OpenStudy (anonymous):

from 3x-3=3 we get x=(3+3)/3 ie. x=2 so 4x(2x+12)= 4 X 2(2 X 2+12)=8 X 16 = 128

OpenStudy (anonymous):

Answer?????????????/ hurry please!!!!!!!!!!!!!!!

OpenStudy (anonymous):

is it 4 times (2x+12) or 4x times(2x+12) 1) 64 2) 128

OpenStudy (anonymous):

-5.997,-.020485,1.01751

OpenStudy (anonymous):

bubbles23 would you clarify the question is it 4 times (2x+12) or is it 4x(2x+12)